OpenStudy (anonymous):

earth is moving around sun& moon is moving around the earth..... what will be the path of the moon , if earth is removed sudenly....help me out?

OpenStudy (anonymous):

as the earth is moving around the sun & moon is moving around the earth so it means the moon and earth both are orbitting the sun together. so if the earth is removed then it will orbit or move around the sun only NOTE: the speed of both the earth and the moon are same while orbiting around sun but as the moon has to orbit around earth too due to which its path teeter a bit so if the Earth suddenly disappear the moon would no longer orbit the place where the Earth was, but it would continue to orbit the Sun.

OpenStudy (anonymous):

Thats a good reply.. to convince urself.. i recommend u to get teh data and calculate what is the orbital speed of the moon around earth and what is the orbital speed of the moon around the sun (which is more or less the same as that of earth) compare the two.. :)

OpenStudy (anonymous):

where will be orbit of moon? why not in the same orbit where the earth revolves around the sun.

OpenStudy (anonymous):

Yes. mv^2/r = GmM/r^2 holds for the sun-moon system as it did for the sun-earth system, once the earth "disappears," and m is the moon mass, M is the sun mass. Depending on the exact geometry [i.e., the phase of the moon], the moon will have approximately the same velocity as did the earth. Then, the relationship between velocity and distance from the earth will be as it was when m represented earth rather than moon, v^2=GM/r.